# Microstructural characterization to reveal evidence of shock deformation in a Campo del Cielo meteorite fragment

### v1

@@ 5 Conclusion
An investigation of isolated Neumann bands using WBV analysis revealed a network of dislocations extending outward from an isolated twin tip and spots of high dislocation density along one side of a Neumann band. The high dislocation density on a single twin boundary was believed to be due to a growth-controlled propagation mechanism dominating for this grain. The high dislocation density and elevated strain energies were found to correspond to incoherent twin boundary side of the Neumann band and may be a result of non-uniform growth on the incoherent twin boundary. Additionally, networks of dislocations were seen to extend outward from the isolated twin tip and are believed to be indicative of further twin propagation through a nucleation-controlled growth mechanism.
Closer investigation was performed into a high strain region of the isolated Neumann bands where a twin-twin intersection was discovered to be the source. Unusual EBSPs were found at the exit point of this intersection with a high intensity {110} band with blurry and lower intensity bands being mirrored across the {110} band. It is suggested this phenomenon was a result of multilayer twinning at the exit point leading to constructive interference for the {110} band and constraint of the crystal structure to the {110} plane due to dislocation slip through the twin plane of the larger Neumann band.
@@
To summarize, EBSD-based investigation of high strain Neumann band regions in the Campo del Cielo meteorite fragment revealed two highly unusual phenomena, the ‘ripple’ microbands across the surface and the mirrored high intensity {110} band EBSP present at a Neumann band intersection. The presence of such phenomena may be valuable in furthering our understanding of high strain rate microstructural changes within both terrestrial and interstellar iron-based materials.
@@ Acknowledgements
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [Discovery grant: RGPIN-2022–04762, ‘Advances in Data Driven Quantitative Materials Characterisation’]. The AMBER-X pFIB-SEM was funded by the BCKDF and CFI-IF (#39798: AM+).

### v2

@@ 5 Conclusion
An investigation of isolated Neumann bands using WBV analysis revealed a network of dislocations extending outward from an isolated twin tip and spots of high dislocation density along one side of a Neumann band. The high dislocation density and elevated strain energies on a single twin boundary were believed to be a result of twin boundary migration through a stepwise growth mechanism or incipient twin intergrowth. Additionally, networks of dislocations were seen to extend outward from the isolated twin tip and are believed to be indicative of further twin propagation through nucleation.
Closer investigation was performed into a high strain region of the isolated Neumann bands where a twin-twin intersection was discovered to be the source. Unusual EBSPs were found at the exit point of this intersection with a high intensity {110} band with blurry and lower intensity bands being mirrored across the {110} band. It is suggested this phenomenon is a result of constrained and localized rotation of the crystal structure about the {110} plane due to dislocation slip associated with the Neumann band-based twin-twin interactions.
@@
To summarize, EBSD-based investigation of high strain Neumann band regions in the Campo del Cielo meteorite fragment revealed two highly unusual phenomena, the ‘ripple’ microbands across the surface and the mirrored high intensity {110} band EBSP present at a Neumann band intersection. This study also reports the coexistence of Neumann bands (microtwins) and microbands within an iron meteorite. These findings may be significant in furthering our understanding of high strain rate microstructural changes within both terrestrial and interstellar iron-based BCC materials under extreme shock conditions.
@@ Acknowledgements
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [Discovery grant: RGPIN-2022–04762, ‘Advances in Data Driven Quantitative Materials Characterisation’]. The AMBER-X pFIB-SEM was funded by the B.C. Knowledge Development Fund (BCKDF) and Canadian Foundation for Innovation-Innovation Fund (CFI-IF) (#39798: AM+). We also acknowledge Haroon Qaiser and Tianbi Zhang for their helpful discussions in creating this work.
CRediT
Graeme J. Francolini: Conceptualization, Formal Analysis, Investigation, Methodology, Writing - Original draft, Writing - review & editing. Thomas. B. Britton Formal Analysis, Funding acquisition, Supervision, Writing - review & editing.
Data Availability
Data will be released to Zenodo when the article is accepted for publication, and a DOI will be inserted here.